If the demand curve for a product is P = 100-Q

Type your answer in the box. 


If the demand curve for a product is P = 100-Q (with P representing price measured in dollars) and the supply curve for the product is P = 4Q, the market clearing price is $_______  value) and the equilibrium quantity is _______ units.
Answer #1

Given

Demand curve P = 100 - QD

supply curve P = 4QS

Therefore, Q D= 100 - P

And Q S= P/4

Equate, these two we get

100 - P = P /4

400 - 4P = P

4P + P = 400

5P = 400

P* = $ 80 per unit

Market clearing price, P* = $ 80 per unit

Plug in P = $ 80 in any equation we get

Q = 100 - P = 100 - 80 = 20 UNITS

Market clearing price $ 80 per unit.

Equilibrium Quantity = 20 units.

Please contact if having any query thank you.